OpenStudy (anonymous):

sinA+sin2A+sin3A+....+sin(nA)?

OpenStudy (anonymous):

Write it backwards and add them together....

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

n/2 ( Sin A + Sin nA)

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

Write it backwards and add them together....

OpenStudy (anonymous):

Ok, thanks.

OpenStudy (anonymous):

i do not think so ,sina +sin(na) is not equal to sin2a+sin{( n-1)a} can not calculate that way

OpenStudy (anonymous):

So far I tried to generalize using the identities: sin (x + b) = sinxcosb + sinbcosx cos (x +b) = cosxcosb - sinxsinb And I got: sinA+sin2A+sin3A+....+sin(​nA) sin2A = sinAcosA + sinAcosA sin2A = 2 sinAcosA sin3A = (2A + A) = sin2A cosA + sinAcos2A 2 sinAcosA*cosA + sinAcos2A 2 sinAcosA*cosA + sinAcos(A+A) 2 sinAcosA*cosA + sinA ( cosA cosA - sinA sinA) 2 sinAcosA*cosA + sinA ( cos^2 A - sin^2A) 2 sinAcos^2 A+ sinA ( cos^2 A - sin^2A) sin^2 A + cos^2 A = 1 sin^2 A = 1 - cos^2 A 2 sinAcos^2 A+ sinA ( cos^2 A - ( 1 - cos^2 A) ) 2 sinAcos^2 A+ sinA ( cos^2 A -1 + cos^2 A ) 2 sinAcos^2 A+ sinA ( 2 cos^2 A -1 ) 2 sinAcos^2 A+ 2 sinA cos^2 A - sinA 4 sin A cos^2 A - sin A sin2A = 2 sinAcosA sin3A = 4 sin A cos^2 A - sin A sin 4A = sin (3A + A) = sin 3A cos A + sin A cos 3A (4 sin A cos^2 A - sin A) cos A + sin A cos 3A (4 sin A cos^2 A - sin A) cos A + sin A cos (2A + A) (4 sin A cos^2 A - sin A) cos A + sin A ( cos 2A cos A - sin 2A sin A) (4 sin A cos^2 A - sin A) cos A + sin A ( (2 cos^2 A -1) cos A - sin 2A sin A)

OpenStudy (anonymous):

‎(4 sin A cos^2 A - sin A) cos A + sin A ( (2 cos^2 A -1) cos A - (2 sinAcosA) sin A) 4 sin A cos^3 A - sin A cos A + 2 sin A cos^3 A -sin A cos A - 2sin^3A cosA 6 sin A cos^3 A - 2 sin A cos A - 2sin^3A cosA 2 sin A cos A ( 3 cos^2 A - 1 - sin^2 A) 2 sin A cos A ( 3 cos^2 A - 1 - (1 - cos^2 A)) 2 sin A cos A ( 3 cos^2 A - 1 - 1 + cos^2 A) 2 sin A cos A ( 4 cos^2 A - 2) 8 sin A cos^3 A - 4 sin A cos A sin 4A = 8 sin A cos^3 A - 4 sin A cos A sin A = sin A sin 2A = 2 sinAcosA sin 3A = 4 sin A cos^2 A - sin A sin 4A = 8 sin A cos^3 A - 4 sin A cos A

OpenStudy (anonymous):

I just thought I made a booboo here.....:-(

OpenStudy (anonymous):

sin A + sin 2A = sin A (1 + 2 cos A) sin A + sin 2 A + sin 3A = sin A + 2 sinA cos A + 4 sin A cos^2 A - sin A sin A ( 1 + 2 cosA + 4 cos^2 A - 1) = sin A (2 cosA + 4cos^2 A) = 2 sinA cos A (1 + 2 cos A) sin A + sin 2A + sin3A + sin 4A = sinA cos A (2 + 4 cos A) + 8 sin A cos^3 A - 4 sin A cos A 2 sinA cos A (1 + 2 cos A) + 2 sin A cos A ( 4 cos^2 A - 2) (2 sinAcosA) ( 1 + 2cosA + 4 cos^2 A -2) (2 sinAcosA) ( 2cosA + 4 cos^2 A -1)

OpenStudy (anonymous):

maybe we need the taylor formula

OpenStudy (anonymous):

I think I'll give up. ://

OpenStudy (anonymous):

No, no...have to split in odds and evens...

OpenStudy (anonymous):

:o

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

Give me a while, not a machine..:-)

OpenStudy (anonymous):

Alright.

OpenStudy (anonymous):

i think this question must belong to infinite series (calcuslu ii)

OpenStudy (anonymous):

Ok, u have a look at that, maybe ur right (but this is finite) For the even part (2A,4A etc) I have n(n+2)/2 SinA CosA (somebody check). Let me see if I can get something out of the odd parts

OpenStudy (anonymous):

This is a right pain, you have 2 cases (n even or odd) The even components are OK (above) but the odd are more trouble. In the end I decided to see what Wolfram had to say, I got Sin(nA/2) Sin ((n+1)A/2) / Sin(A/2) I think there is some way to do this with complex numbers, not sure, though...

OpenStudy (anonymous):

Actually, maybe that should simplify to just Sin ((n+1)A/2) I see the way to deal with the odd even u need half angle so multiply the whole shebang by sin (A/2) and use sin x sin y = 1/2 (cos(x-y) -cos(x+y)) and then it just frops out:-):-)

OpenStudy (anonymous):

Drops out I mean...

OpenStudy (anonymous):

Because all the half angles just cancel each other out...

OpenStudy (zarkon):

\[\sum_{k=1}^{n}\sin(k A) = \frac{\sin\left(\frac{n A}{2}\right) \sin\left(\frac{ (n+1) A}{2}\right)}{\sin\left(\frac{A}{2}\right)}\] Can be proven by indiction

OpenStudy (anonymous):

Yes, that's same as Wolfram, u don't need to do that....

OpenStudy (anonymous):

okay I got this \[SinA + Sin B = 2Sin \frac{A+B}{2} Cos\frac{A -B}{2}\] If we do this with odd terms we get even sin terms

OpenStudy (anonymous):

True, but the way I just said is dead easy, about 3 or 4 lines is all and u get Sin ((n+1)A/2)

OpenStudy (anonymous):

Oh yeah sorry i didn't read that just started replying

OpenStudy (anonymous):

Well estudier you talk about complex number way of solving it

OpenStudy (anonymous):

not talk about but mentioned it

OpenStudy (anonymous):

what if we take sin a to cos (pi/2 - a)

OpenStudy (anonymous):

then imaginary is zero

OpenStudy (anonymous):

and pi/2 -a = npi

OpenStudy (anonymous):

then the whole angle becomes pi x n ..then maybe polar form e^i pi n now n should be integers

OpenStudy (anonymous):

ah i don't know anyway your method is better thanks for that

OpenStudy (anonymous):

Sorry, phone call...I will have a look at that...

OpenStudy (anonymous):

: ) thanks

OpenStudy (anonymous):

U can use the complex equivalent of half angle set up. If you multiply series by i and add it to Cos A + Cos 2A etc u get a complex sum of form a + ib then use Euler formula to get cos A = 1/2( e^iA + e^iA) and sin x = 1/2 (e^iA - e^iA) -> e^iA + e^2iA + etc which is geometric series and then you just compare real and imaginary terms it comes out to the Wolfram answer I put up before (I made a silly cancelling error in the trig way I did it). Still think the trig way is more straightforward once u see the half angle trick.